# Do Language Models Understand the Cognitive Tasks Given to Them? Investigations with the N-Back Paradigm

### v1

@@ Abstract
Cognitive tasks originally developed for humans are now increasingly used to study language models. While applying these tasks is often straightforward, interpreting their results can be challenging. In particular, when a model underperforms, it’s often unclear whether this results from a limitation in the cognitive ability being tested or a failure to understand the task itself. A recent study argued that GPT 3.5’s declining performance on 2-back and 3-back tasks reflects a working memory capacity limit similar to humans (Gong et al., 2024). By analyzing a range of open-source language models of varying performance levels on these tasks, we show that the poor performance instead reflects a limitation in task comprehension and task set maintenance. In addition, we push the best performing model to higher n values and experiment with alternative prompting strategies, before analyzing model attentions. Our larger aim is to contribute to the ongoing conversation around refining methodologies for the cognitive evaluation of language models.
@@ 1 Introduction
A recent study applied the n-back task (Figure 1) to GPT 3.5 and concluded from the model’s poor 2-back and 3-back performance that it had a working memory capacity limit of approximately 3, apparently similar to humans (Gong et al., 2024). This interpretation raises two concerns. First, while working memory capacity limits are well established in human cognition, we cannot assume these same constraints exist or can be meaningfully measured in language models. Second, these results may simply reflect the model’s failure to understand the task requirements rather than any inherent memory limitation.
@@
The remainder of this paper is organized as follows. Section 2 covers relevant background. Section 3 introduces the dataset, prompting approach, and evaluation metrics. In Section 4.1, we evaluate a range of language models on the n-back task, focusing on the letters retrieved for comparison with the current ones. In Section 4.2, we show that, while successful models correctly map n-back instructions to n-back consistent retrievals (i.e., assign the highest log probabilities to n-back consistent retrievals when given n-back instructions), weaker models would map 2-back and 3-back instructions to 1-back consistent retrievals. We also analyze the effect of task instructions and demonstrations on this mapping. In Section 4.3, we find that, while the intermediate models are initially able to follow 2-back and 3-back instructions, as more stimuli come in, they tend to drift towards 1-back and 2-back consistent retrievals, respectively. In Section 4.4, we challenge the best model to perform 1-back through 10-back tasks and notice a signature of task comprehension: the model assigns the highest log probabilities to correct n-back continuations while assigning progressively lower log probabilities to m-back continuations as m deviates from n. In Sections 4.5 to 4.7, we discuss additional experiments with alternative prompting strategies for comparison. In Section 4.8, we identify an attention pattern whose prevalence predicts 2-back task performance.
@@ 2 Background and Related Work
Such evaluations are susceptible to both overclaiming and underclaiming. On the one hand, false positives can result from training data contamination (Sainz et al., 2023), potentially compromising the validity of vignette-based assessments where models may produce memorized responses. On the other hand, underestimation of model capabilities can happen when we erroneously assume task comprehension, especially for smaller models (Hu and Frank, 2024). Prior studies have also investigated how well language models adhere to prompt instructions (Webson and Pavlick, 2022) and how their performance compares to that of humans (Webson et al., 2023). In light of other methodological challenges in cognitive evaluations of language models, such as prompt sensitivity and cultural biases, Ivanova 2023 outlines recommendations for best practices.
@@
One of the most salient and mysterious aspects of human working memory is its severely constrained capacity (Miller, 1956; Cowan, 2012). This capacity limit is often measured using the n-back task (Kirchner, 1958). To the best of our knowledge, Gong et al. 2024 was the first to apply the n-back task to a language model, specifically the GPT 3.5 Turbo variant of ChatGPT. They experimented with different prompting strategies, including those incorporating feedback and reasoning. As n increased from 1 to 3, they observed a sharp decline in model performance and concluded that ChatGPT had a working memory capacity limit of approximately 3.
@@ 3.1 Data and Prompts
$\texttt{SYS: }&\texttt{[TASK INSTRUCTION]}\\[-2.41109pt] \texttt{USR: }&\texttt{k}\\[-2.41109pt] \texttt{LLM: }&\texttt{k and none are different.}\\[-2.41109pt] \texttt{USR: }&\texttt{k}\\[-2.41109pt] \texttt{LLM: }&\texttt{k and k are identical.}\\[-2.41109pt] \texttt{USR: }&\texttt{a}\\[-2.41109pt] \texttt{LLM: }&\texttt{a and k are different.}\\[-2.41109pt] &\vdots\end{aligned}\right\}\texttt{DEMO}\\[3.44444pt] &\left.\begin{aligned} \texttt{SYS: }&\texttt{[TASK INSTRUCTION]}\\[-2.41109pt] \texttt{USR: }&\texttt{e}\\[-2.41109pt] \texttt{LLM: }&\underline{\texttt{e and none are different.}}\\[-2.41109pt] \texttt{USR: }&\texttt{f}\\[-2.41109pt] \texttt{LLM: }&\underline{\texttt{f and e are different.}}\\[-2.41109pt] \texttt{USR: }&\texttt{f}\\[-2.41109pt] \texttt{LLM: }&\underline{\texttt{f and f are identical.}}\\ &\vdots\end{aligned}\right\}\texttt{TEST}\end{aligned}$
@@ 3.2 Models
We use GPT 3.5 Turbo and open-source instruction-tuned models from the Qwen (Bai et al., 2023), Llama (Dubey et al., 2024), and Gemma (Team et al., 2024) families. Each model is prompted recursively to complete the tasks. For the open-source models, we analyze the token log probabilities and attention patterns in addition to the generated responses.
@@ 3.3 Metrics
But how can we be sure that a model has inferred the right task from the instruction? One hypothesis is that, despite being prompted to do the n-back task, the models might be following m-back instructions instead. To investigate this, we adopt counterfactual measures by providing n-back instructions and evaluating the accuracies and log probabilities of retrievals consistent with the m-back task. We also apply variants of these measures, which we will detail in later sections.
@@ 4.1 Task Performance
  • T3 models achieve nearly perfect retrieval accuracy on 1-back trials, but their performance drops to around 20% or lower on 2-back and 3-back trials.
  • T2 models achieve nearly perfect retrieval accuracy on 1-back trials and around 50% and 40% on 2-back and 3-back trials, respectively.
  • T1 models achieve 100% retrieval accuracy on 1-back trials and over 80% on 2-back and 3-back trials.
@@
For subsequent analyses, we select the best performing model, Llama 3.1 70b Instruct (T1), the worst performing model, Qwen 1.5 14b Chat (T3), and Gemma 2 27b Instruct (T2) to represent each performance tier.
@@ 4.2 Task Comprehension
Under 1-back instructions, $\mathsf{P}_{1,1}>\mathsf{P}_{1,2}>\mathsf{P}_{1,3}$ across all models. The same is true when no task demonstrations are provided, with no significant difference between $\mathsf{P}_{1,m}$ and $\mathsf{P}^{-}_{1,m}$ for $m=1,2,3,$ as shown in Figure 3. Overall, this is unsurprising, given the near-perfect performance of all models on the 1-back trials.
@@ 4.3 Task Set Maintenance
Each n-back trial presents a sequence of 24 letters. Successful task completion requires consistent adherence to the task instruction as more stimuli are presented. Here, we investigate whether language models show a progressive decline in their ability to produce n-back consistent responses over time. Previously, performance metrics were averaged across time steps for each trial; now, we average across trials for each time step.
Specifically, at each time step $i$ in the n-back task, we measure the average accuracy of m-back consistent retrievals for each $m\leq n$, given the model’s own responses up to time step $i-1$. Denote the accuracy as $\mathsf{A}_{n,\cdot}(i,m)$.
@@
Despite 2-back instructions and demonstrations, the T2 model gradually drifts towards 1-back consistent responses as the task progresses, suggesting that the accumulation of 1-back consistent errors may have significantly biased subsequent responses. To test this hypothesis, we manipulate the model’s response history by providing m-back consistent responses for i steps following n-back instructions and demonstrations, then compute the average m-back accuracy for time steps $i+1$ through 24, which we denote as $\mathsf{A}_{n,m}(i+1:24,m)$. Figure 8 shows that, as 1-back errors accumulate, 1-back responses are increasingly favored by the T2 model for subsequent steps, despite 2/3-back instructions and demonstrations. In comparison, both $\mathsf{A}_{2,2}(i+1:24,2)$ and $\mathsf{A}_{3,3}(i+1:24,3)$ remain relatively low, showing that correct responses do not bias subsequent answers to the same degree.
Figure 8: Top: $\mathsf{A}_{2,m}(i+1:24,m)$ for $m=1,2,$ and $3\leq i\leq 23$, using Gemma 2 27b Instruct (T2). Bottom: $\mathsf{A}_{3,m}(i+1:24,m)$ for $m=1,2,3,$ and $4\leq i\leq 23$, using the same model.
Figure 9: 1-back to 10-back accuracies for Llama 3.1 70b Instruct with (bottom) and without (top) curriculum learning. Each full bar corresponds to the task (identical/different categorization) accuracy. The blue portion corresponds to retrieval accuracy.
@@ 4.4 T1 Model Performance as N Increases
Figure 9 shows that the retrieval accuracy gradually declines as n increases, although, even at $n=8,9,10,$ the model is still able to exactly retrieve the correct letters 75.25%, 66.08% and 57.1% of the times, which translates to task accuracies of 83.33%, 78.25% and 71.92%.
@@ 4.5 Curriculum Learning
The practice of training models on examples of increasing difficulty is known in machine learning as curriculum learning (Bengio et al., 2009). Here, we repeat the experiments from Section 4.4 with in-context curriculum learning to gradually familiarize the model with the task. Specifically, before prompting Llama 3.1 70b Instruct to perform an n-back task, we provide demonstrations that include letter sequences and corresponding correct responses for tasks ranging from 1-back to n-back. As shown in Figure 9, this approach leads to significant improvements in performance for larger n values. The model achieves retrieval accuracies of 79.83%, 80.17% and 71.67% and task accuracies of 90.08%, 90.08% and 84.75%, for $n=8,9,10$.
@@ 4.8 Attention Analysis
Attentions in transformer-based language models reveal how much each generated token attends to every preceding token. We hypothesize that, for each retrieval, a more performant model should attend more to the source token from n steps back. This is precisely what we observe in the Qwen models. For each (trial, layer, head), we obtain the mean retrieval attention (MRAT) by averaging the attention each retrieval token gives to the correct source token. Indeed, compared to the 14b model, Qwen 2 72b Instruct (T1) shows a higher concentration at the tail end of the MRAT distribution (Figure 12). Figure 13 shows the attention map with the highest MRAT, which almost exactly matches our hypothesized pattern. However, Llama models do not exhibit this pattern to the same degree. Attentions in Llama 3.1 models are much more diffuse; The maximum MRATs for Llama 3.1 8b Instruct and Llama 3.1 70b Instruct are 4.86% and 8.52%, respectively.
@@ 5 Conclusion
In this work, we apply the n-back task, a common working memory test, to a range of language models. We categorize the models into three performance tiers based on retrieval accuracy. Through analyzing a representative model from each tier, we identify three distinct levels of task understanding and task set maintenance capabilities, highlighting their confounding nature. We challenge the best model to perform 1 through 10-back tasks, noticing a gradual decline in performance and the benefit of in-context curriculum learning for larger n’s. We find that interactive demos, though closer to human study paradigms, are less effective at conveying the task. For comparison, we also include results under an alternative answer format that allows for rehearsal. Finally, we observe that more performant models tend to have higher retrieval attentions, although the values for Llama models seem surprisingly small.
@@ 6 Limitations
Despite our careful selection of prompts and experimentation with various prompting strategies, the potential for more effective prompts or techniques to enhance task understanding remains. Future research could investigate systematic approaches to explore a wider range of prompts and their impact on task understanding and task set maintenance.
@@
Another limitation is that we do not examine the internal model circuits that may be responsible for inferring and maintaining task sets. However, our experiments with the n-back paradigm provides a good starting point for future research. Causal interventions on smaller models may yield insights into the underlying mechanisms.

### v2

@@ Abstract
Cognitive tasks originally developed for humans are now increasingly used to study language models. While applying these tasks is often straightforward, interpreting their results can be challenging. In particular, when a model underperforms, it is often unclear whether this results from a limitation in the cognitive ability being tested or a failure to understand the task itself. A recent study argues that GPT 3.5’s declining performance on 2-back and 3-back tasks reflects a working memory capacity limit similar to humans (Gong et al., 2024). By analyzing a range of open-source language models of varying performance levels on these tasks, we show that the poor performance instead reflects a limitation in task comprehension and task set maintenance. In addition, we challenge the best-performing model with progressively harder versions of the task (up to 10-back) and experiment with alternative prompting strategies, before analyzing model attentions. Our larger aim is to contribute to the ongoing conversation around refining methodologies for the cognitive evaluation of language models.
@@ 1 Introduction
A recent study applies the n-back task (Figure 1) to GPT 3.5 and concludes from the model’s poor 2-back and 3-back performance that it has a working memory capacity limit of approximately 3, apparently similar to humans (Gong et al., 2024). This interpretation raises two concerns. First, while working memory capacity limits are well established in human cognition, we cannot assume these same constraints exist or can be meaningfully measured in language models. Second, these results may simply reflect the model’s failure to understand the task requirements rather than any inherent memory limitation.
@@
The remainder of this paper is organized as follows. Section 2 covers relevant background. Section 3 introduces the dataset, prompting approach, and evaluation metrics. In Section 4.1, we evaluate a range of language models on the n-back task, focusing on the stimuli retrieved for comparison with the current ones. In Section 4.2, we show that, while successful models correctly map n-back instructions to n-back consistent retrievals (i.e., assign the highest log probabilities to n-back consistent retrievals when given n-back instructions), weaker models would map 2-back and 3-back instructions to 1-back consistent retrievals. We also analyze the effect of task instructions and demonstrations on these mappings. In Section 4.3, we find that, while the intermediate models are initially able to follow 2-back and 3-back instructions, as more stimuli come in, they tend to drift toward 1-back and 2-back consistent retrievals, respectively. In Section 4.4, we challenge the best model to perform 1-back through 10-back tasks and notice a signature of task comprehension: the model assigns the highest log probabilities to correct n-back continuations while assigning progressively lower log probabilities to m-back continuations as m deviates from n. In Sections 4.5 to 4.7, we discuss additional experiments with alternative prompting strategies for comparison. In Section 4.8, we identify an attention pattern whose prevalence predicts 2-back task performance.
@@ 2 Background and Related Work
Such evaluations are susceptible to both overclaiming and underclaiming. On the one hand, false positives can result from training data contamination (Sainz et al., 2023), potentially compromising the validity of vignette-based assessments where models may produce memorized responses. On the other hand, underestimation of model capabilities can happen when we erroneously assume task comprehension, especially for smaller models (Hu and Frank, 2024). Prior studies have also investigated how well language models adhere to prompt instructions (Webson and Pavlick, 2022) and how their performance compares to that of humans (Webson et al., 2023). In light of other methodological challenges in the cognitive evaluation of language models, such as prompt sensitivity and cultural biases, Ivanova 2023 outlines recommendations for best practices.
@@
One of the most salient and mysterious aspects of human working memory is its severely constrained capacity (Miller, 1956; Cowan, 2012). This capacity limit is often measured using the n-back task (Kirchner, 1958). To the best of our knowledge, Gong et al. 2024 is the first to apply the n-back task to a language model, specifically the GPT 3.5 Turbo variant of ChatGPT. They experiment with different prompting strategies, including those incorporating feedback and reasoning. As n increases from 1 to 3, they observe a sharp decline in model performance and conclude that ChatGPT has a working memory capacity limit of approximately 3.
@@ 3.1 Data and Prompts
$\texttt{SYS: }&\texttt{[TASK INSTRUCTIONS]}\\[-2.41109pt] \texttt{USR: }&\texttt{k}\\[-2.41109pt] \texttt{LLM: }&\texttt{k and none are different.}\\[-2.41109pt] \texttt{USR: }&\texttt{k}\\[-2.41109pt] \texttt{LLM: }&\texttt{k and k are identical.}\\[-2.41109pt] \texttt{USR: }&\texttt{a}\\[-2.41109pt] \texttt{LLM: }&\texttt{a and k are different.}\\[-2.41109pt] &\vdots\end{aligned}\right\}\texttt{DEMO}\\[3.44444pt] &\left.\begin{aligned} \texttt{SYS: }&\texttt{[TASK INSTRUCTIONS]}\\[-2.41109pt] \texttt{USR: }&\texttt{e}\\[-2.41109pt] \texttt{LLM: }&\underline{\texttt{e and none are different.}}\\[-2.41109pt] \texttt{USR: }&\texttt{f}\\[-2.41109pt] \texttt{LLM: }&\underline{\texttt{f and e are different.}}\\[-2.41109pt] \texttt{USR: }&\texttt{f}\\[-2.41109pt] \texttt{LLM: }&\underline{\texttt{f and f are identical.}}\\ &\vdots\end{aligned}\right\}\texttt{TEST}\end{aligned}$
@@ 3.2 Models
We use GPT 3.5 Turbo and open-source instruction-tuned models from the Qwen (Bai et al., 2023), Llama (Dubey et al., 2024), and Gemma (Team et al., 2024) families. Each model is prompted recursively to complete the trials. For the open-source models, we analyze the token log probabilities and attention patterns in addition to the generated responses.
@@ 3.3 Metrics
But how can we be sure that a model has inferred the right task from the instructions? One hypothesis is that, despite being prompted to do the n-back task, the model might be following m-back instructions instead. To investigate this, we adopt counterfactual measures by providing n-back instructions and evaluating the accuracies and log probabilities of retrievals consistent with the m-back task. We also apply variants of these measures, which we detail in later sections.
@@ 4.1 Task Performance
  • T3 models achieve nearly perfect retrieval accuracies on 1-back trials, but their performance drops to around 20% or lower on 2-back and 3-back trials.
  • T2 models achieve nearly perfect retrieval accuracies on 1-back trials and around 50% and 40% on 2-back and 3-back trials, respectively.
  • T1 models achieve 100% retrieval accuracies on 1-back trials and over 80% on 2-back and 3-back trials.
@@
For subsequent analyses, we select the best-performing model, Llama 3.1 70b Instruct (T1), the worst-performing model, Qwen 1.5 14b Chat (T3), and Gemma 2 27b Instruct (T2) to represent each performance tier.
@@ 4.2 Task Comprehension
Under 1-back instructions, $\mathsf{P}_{1,1}>\mathsf{P}_{1,2}>\mathsf{P}_{1,3}$ across all models. The same is true when no task demonstrations are provided, with no significant difference between $\mathsf{P}_{1,m}$ and $\mathsf{P}^{-}_{1,m}$ for $m=1,2,3,$ as shown in Figure 3. Overall, this is unsurprising, given the near-perfect performance of all models on 1-back trials.
@@ 4.3 Task Set Maintenance
Each n-back trial presents a sequence of 24 letters. Successful task completion requires consistent adherence to the task instructions as more stimuli are presented. Here, we investigate whether language models show a progressive decline in their ability to produce n-back consistent responses over time. Previously, performance metrics were averaged across time steps for each trial; now, we average across trials for each time step.
Specifically, at each time step $i$ in the n-back task, we measure the average accuracy of m-back consistent retrievals for each $m\leq n$, given the model’s own responses up to time step $i-1$. Denote this as $\mathsf{A}_{n,\cdot}(i,m)$.
@@
Despite 2-back instructions and demonstrations, the T2 model gradually drifts toward 1-back consistent responses over time, suggesting that the accumulation of 1-back consistent errors may have significantly biased subsequent responses. To test this hypothesis, we manipulate the model’s response history by providing m-back consistent responses for i steps following n-back instructions and demonstrations, then compute the average m-back accuracy for time steps $i+1$ through 24, which we denote as $\mathsf{A}_{n,m}(i+1:24,m)$. Figure 8 shows that, as 1-back errors accumulate, 1-back responses are increasingly favored by the T2 model for subsequent steps, despite 2/3-back instructions and demonstrations. In comparison, both $\mathsf{A}_{2,2}(i+1:24,2)$ and $\mathsf{A}_{3,3}(i+1:24,3)$ remain relatively low, showing that correct responses do not bias subsequent answers to the same degree.
Figure 8: Top: $\mathsf{A}_{2,m}(i+1:24,m)$ for $m=1,2$ and $3\leq i\leq 23$, using Gemma 2 27b Instruct (T2). Bottom: $\mathsf{A}_{3,m}(i+1:24,m)$ for $m=1,2,3$ and $4\leq i\leq 23$, using the same model.
Figure 9: 1-back to 10-back accuracies for Llama 3.1 70b Instruct with (bottom) and without (top) curriculum learning. Each full bar corresponds to task (identical/different categorization) accuracy. The blue portion corresponds to retrieval accuracy.
@@ 4.4 T1 Model Performance as N Increases
Figure 9 shows that the retrieval accuracy gradually declines as n increases; although, even at $n=8,9,10,$ the model is still able to exactly retrieve the correct letters 75.25%, 66.08%, and 57.1% of the times, which translates to task accuracies of 83.33%, 78.25%, and 71.92%.
@@ 4.5 Curriculum Learning
The practice of training models on examples of increasing difficulty is known in machine learning as curriculum learning (Bengio et al., 2009). Here, we repeat the experiments from Section 4.4 with in-context curriculum learning to gradually familiarize the model with the task. Specifically, before prompting Llama 3.1 70b Instruct to perform an n-back task, we provide demonstrations that include letter sequences and corresponding correct responses for tasks ranging from 1-back to n-back. As shown in Figure 9, this approach leads to significant improvements in performance for larger n values. The model achieves retrieval accuracies of 79.83%, 80.17%, and 71.67% and task accuracies of 90.08%, 90.08%, and 84.75%, for $n=8,9,10$.
@@ 4.8 Attention Analysis
Attentions in transformer-based language models reveal how much each generated token attends to every preceding token. We hypothesize that, for each retrieval, a more performant model should attend more to the source token from n steps back. This is precisely what we observe in the Qwen models. For each (trial, layer, head), we obtain the mean retrieval attention (MRAT) by averaging the attention each retrieval token gives to the correct source token. Indeed, compared to the 14b model, Qwen 2 72b Instruct (T1) shows a higher concentration at the tail end of the MRAT distribution (Figure 12). Figure 13 shows the attention map with the highest MRAT, which almost exactly matches our hypothesized pattern. However, Llama models do not exhibit this pattern to the same degree. Attentions in Llama 3.1 models are much more diffuse. The maximum MRATs for Llama 3.1 8b Instruct and Llama 3.1 70b Instruct are 4.86% and 8.52%, respectively.
@@ 5 Conclusion
In this work, we apply the n-back task, a common working memory test, to a range of language models. We categorize the models into three performance tiers based on retrieval accuracy. Through analyzing a representative model from each tier, we identify three distinct levels of task understanding and task set maintenance capabilities, highlighting their confounding nature. We challenge the best model to perform 1 through 10-back tasks, noticing a signature of task comprehension and the benefit of in-context curriculum learning for larger n’s. We find that interactive demos, though closer to human study paradigms, are less effective at conveying the task. For comparison, we also include results under an alternative answer format that allows for rehearsal. Finally, we observe that more performant models tend to have higher retrieval attentions, although the values for Llama models seem surprisingly small.
@@ 6 Limitations
Despite our careful selection of prompts and experimentation with various prompting strategies, the potential for more effective prompts or techniques to enhance task understanding remains.
@@
Another limitation is that we do not examine the internal model circuits that may be responsible for inferring and maintaining task sets. However, our experiments with the n-back paradigm provide a good starting point for future research. Causal interventions on smaller models may yield insights into the underlying mechanisms.